OpenStudy (anonymous):

HELP ME PLEASE. I'LL GIVE A MEDAL!! A 1.57 g sample of fencing wire consists mostly of iron. the sample was dissolved in sulfuric acid and the resulting solution made up to exactly 250 ml. this solution reacted completely with 270 ml of 0.020 mol/L KMnO4 solution. Determine the percentage of iron in the wire. Mno4 + 5Fe + 8H >> Mn + 5Fe + 4H2O

OpenStudy (aaronq):

Find the moles of iron from the titration. To do so, first find the moles of KMnO4 used with: \(\sf Molarity=\dfrac{moles}{Molar~mass}\) After use the coefficients in the balanced reaction and set up a ratio using the species of interest, like so: e.g. for a general reaction: \(\sf \color{red}{a}A + \color{blue}{b}B\) \(\rightleftharpoons\) \( \color{green}{c}C\) where upper case are the species (A,B,C), and lower case (a,b,c) are the coefficients , \(\sf \dfrac{n_A}{\color{red}{a}}=\dfrac{n_B}{\color{blue}{b}}=\dfrac{n_C}{\color{green}{c}}\) From here you can isolate what you need. For example: if you have 2 moles of B, how many moles of C can you produce? solve algebraically: \(\sf\dfrac{2~mol}{\color{blue}{b}}=\dfrac{n_C}{\color{green}{c}}\rightarrow n_C=\dfrac{2*\color{green}{c}}{\color{blue}{b}}\) For your example, your ratio is: \(\large \dfrac{n_{MnO_4^-}}{1}=\dfrac{n_{Fe}}{5}\) After convert moles to mass and find the percentage of Fe in the sample.

OpenStudy (aaronq):

I wrote the wrong formula for molarity, it should be: \(\sf Molarity=\dfrac{moles~of~solute}{L_{solution}}\)